OpenStudy (anonymous):

Find The equation of the tangent to the curve y=x^2+x^3 at the point (1,0) .

OpenStudy (ash2326):

@Eyad We have to find the equation of the tangent to the curve \(y=x^2+x^3\) at the point (1, 0) First find \(\frac{dy}{dx}\) at the point (1, 0) this will be the slope of the tangent at (1, 0) and then you can find the equation using this \[\frac{(y-y1)}{(x-x1)}=m\] where \[(x1,y1)=(1, 0) \] and \[\large m= ( \frac{dy}{dx})_{(1,0)}\] Can you try now?

OpenStudy (anonymous):

so its gonna be \[\frac{y-0}{x-1}=0\] ?

OpenStudy (anonymous):

@ash2326 .

OpenStudy (ash2326):

\[\frac{dy}{dx}=3x^2+2x\] put x=1 and y=0 \[\frac{dy}{dx}=3+2=5\]

OpenStudy (anonymous):

Oh Sry i think i need some sleep xD, U did great :D Although its \[2x-3x^2 ...\] where the slope is -1 TY ^_^

OpenStudy (ash2326):

You really are sleepy:P you wrote y=x^2+x^3 instead of y=x^2-x^3

OpenStudy (anonymous):

OH LOL ,Its y=x^2-x^3 .. Night :D